- with B in QIII, find the following. M18032. C w ereldong animolla til uova 34. s 36.
#35 please
Answer #1

Sin B = -% () COMCE) = ? Si come COD B =-|- SinB = 11-1143) Ces () = + (0-1303) Et 3-252 V 3+ 252 = + 13- 2.52342 12 1 3+2 52